OpenStudy (firejay5):

Lunch Combo Meals: 1. Two tacos and 1 burrito = $6.55 2. One Enchilada, 1 taco, and a burrito = $7.10 3. Two enchiladas and two tacos = $8.90 If Maka wants 2 burritos and 1 enchilada, how much should she plan to spend???

OpenStudy (anonymous):

2t + b = 6.55 t + e + b = 7.10 2t + 2e = 8.90 b = 6.55-2t (from 1st equation) e = 4.45 -t (from 3rd equation) Substitute into 2nd equation to get: t + (4.45-t) + (6.55-2t) = 7.10 -2t +11 = 7.10 -2t = -3.90 t = 1.95 b = 2.65

OpenStudy (anonymous):

Dear mod who just warned me on this, You're a joke. He already got the answer and was asking for confirmation. Go find something useful to do with your time. Goodnight.

OpenStudy (firejay5):

What are you talking about @SmoothMath

OpenStudy (anonymous):

Some mod warned me for my answer to this question, saying I should have guided you to the answer instead of solving it for you. Pretty hilarious, since you had already solved the problem on your own.

OpenStudy (firejay5):

You know what @SmoothMath

OpenStudy (anonymous):

But that's representative of the mods on this site. They waste their time nit picking high quality users causing no trouble, and at the same time, they ignore the users who are actually dragging down the quality of the site.

OpenStudy (firejay5):

@SmoothMath guess what

OpenStudy (anonymous):

What?

OpenStudy (firejay5):

It doesn't matter as long as I got the answer and understood it; it's fine

OpenStudy (anonymous):

Right, I know. I still want to call the people out who run this site. They need to be told when they're doing a terrible job.

OpenStudy (omnomnom):

yes .... i think

OpenStudy (anonymous):

She can bye them seperately or must buy a combo to get them.

OpenStudy (anonymous):

@Firejay5

OpenStudy (firejay5):

yes @sauravshakya

OpenStudy (anonymous):

What she can bye them seperately?

OpenStudy (firejay5):

not how she should spend it

OpenStudy (anonymous):

2t + b = 6.55 t + e + b = 7.10 2t + 2e = 8.90 SOLVE that three equations to get the value of t,e and b

OpenStudy (firejay5):

that's not what I am asking

OpenStudy (anonymous):

Then what?

OpenStudy (firejay5):

If Maka wants 2 burritos and 1 enchilada, how much should she plan to spend???

OpenStudy (anonymous):

Find: 2b+e

OpenStudy (firejay5):

Would I take the answers from 27 to do that

OpenStudy (anonymous):

can anyone hep me on some computer science